OpenStudy (anonymous):

Calculate the volume of a sphere with radius measuring 5.4 cm

jhonyy9 (jhonyy9):

V =4/3 (pir^3) - if i remember right

OpenStudy (anonymous):

helppppppppppppppppppppppppppppppppppppppppppppppppppppppppppppppppppppppppppppppppppppppppppppppppppppppppppppppppppppppppppppppppppppppppppppppppp

jhonyy9 (jhonyy9):

- you know formula for volum of sphere ?

OpenStudy (anonymous):

noooooooooooooooooooooo

jhonyy9 (jhonyy9):

- yes is right how i have wrote there

jhonyy9 (jhonyy9):

- you need just make the calcules

OpenStudy (anonymous):

yout dumb

OpenStudy (anonymous):

yout dumb

jhonyy9 (jhonyy9):

V=4/3(pi*r^3) =4/3 (pi*157,464) =209,952 cm^3

jhonyy9 (jhonyy9):

ok ?